40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Mar 7 15:26:10 2021, gautam, Update, LSC, Housekeeping + more PRMI PRMI_SBres.pngMICH_act_calib.pdf
    Reply  Mon Mar 8 22:01:26 2021, gautam, Update, LSC, More PRMI PRMI1f_noArmssensMat.pdfPRMI3f_noArmssensMat.pdf
Message ID: 15875     Entry time: Sun Mar 7 15:26:10 2021     Reply to this: 15883
Author: gautam 
Type: Update 
Category: LSC 
Subject: Housekeeping + more PRMI 
  1. Beam pointing into PMC was tweaked to improve transmission.
  2. AS110 photodiode was re-installed on the AS table - I picked off 30% of the light going to the AS WFS using a beamsplitter and put it on the AS110 photodiode.
  3. Adjusted ASDC whitening gain - we have been running nominally with +18dB, but after Sept 2020 vent, there is ~x3 amount of light incident on the AS55 RFPD (from which the ASDC signal is derived). I want to run the dither alignment servos that use this PD using the same settings as before, hence this adjustment.
  4. Adjusted digital demod phases of POP22, POP110 and AS110 signals with the PRMI locked (sideband resonant). I want these to be useful to debug the PRMI. the phases were adjusted so that AS110_Q, POP22_I and POP110_I contain the signal (= sideband buildup) when the PRMI is locked.
  5. Ran the actuator calibration routine for BS, ITMX and ITMY - i'll try and do the PRM and ETMs as well later.
  6. With the PRMI locked (sidebands resonant), looked at the sideband power buildup. POP22 and POP110 remain stable, but there is some low frequency variation in the AS110_Q channel (but not the I channel, so this is really a time varying transmission of the f2 sideband to the dark port). What's that about? Also unsure about those abrupt jumps in the POP22/POP110 signals, see Attachment #1 (admittedly these are slow channels). I don't see any correlation in the MICH control signal.
  7. Measured the loop shapes of the MICH (UGF ~90 degrees, PM~30 degrees) and PRCL (UGF~110 Hz, PM~30 degreees) loops - stability margins and loop UGFs seem reasonable to me.
  8. Tried nulling the MICH-->PRCL coupling by adjusting the MICH-->PRM matrix element - as has been the case for a while, unable to do any better, and I can't null that line as we expect to be able to.
  9. Not expecting to get anything sensible, but ran some sensing matrix lines (at the correct frequencies this time).
  10. Tried locking the PRMI with MICH actuation to an ITM instead of the BS - I can realize the lock but the loop OLTF I measure with this configuration is very weird, needs more investigation. I may look into this later today evening.

I was also reminded today of the poor reliability of the LSC whitening electronics. Basically, there may be hidden saturations in all the channels that have a large DC value (e.g. the photodiode DC mon channels) due to the poor design of the cascaded gain stages. I was thinking about using the REFL DC channel to estimate the mode-matching into the PRC, but this has a couple of problems. Electronically, there may be some signal distortion due to the aforementioned problem. But in addition, optically, the estimation of mode-matching into the PRC by comparing REFL DC levels in single bounce off the PRM and the PRMI locked has the problem that the mode-matching is degenerate with the intra-cavity loss, which is of the same order as the mode mismatch (a percent or two I claiM). If Koji or someone else can implement the fix suggested by Hartmut for all the LSC whitening channels, that'd give us more faith in the signals. It may be less work than just replacing all the whitening filters with a better design (e.g. the aLIGO ISC whitening filter which implements the cascaded gain stages using single OP27s and more importantly has a 1 kohm series resistance with the input to the op amp (so the preceeding stage never has to drive > 10V/1kohms ~10mA of DC current) would presumably reduce distortion.

Attachment 1: PRMI_SBres.png  40 kB  | Hide | Hide all
PRMI_SBres.png
Attachment 2: MICH_act_calib.pdf  34 kB  Uploaded Sun Mar 7 18:17:44 2021  | Hide | Hide all
MICH_act_calib.pdf
ELOG V3.1.3-